# Estimates for trilinear and quadrilinear character sums

### v3

@@ 6.2. Bounding $\Sigma_{2}$
We substitute the bounds (6.11), (6.12) and (6.21) into (6.10) and derive
$Q_{2}(d)$ $\leqslant p^{o(1)}\sup_{\begin{subarray}{c}C\leqslant K/d\end{subarray}}\left(CL^{2}M^{2}NUWd^{-1}\right)^{1-\frac{1}{r}}(M/C+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\mathscr{L}_{2}^{2}LM(CW+NU/d)d^{-2}V^{-1}\right)^{\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}$
$\leqslant p^{o(1)}\left(KL^{2}M^{2}NUWd^{-2}\right)^{1-\frac{1}{r}}(dM/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\times\left(\mathscr{L}_{2}^{2}KLMNd^{-3}V^{-2}\right)^{\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}.$
Recalling (6.8) and (6.9) and then summing over $d\leqslant D$, we obtain
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+KL^{2}M^{2}U^{-1}p^{o(1)}+KL^{2}M\left(1+\frac{MN}{pV}\right)p^{o(1)}.$
@@
$\qquad\qquad+(LM)^{2}p^{\frac{1}{r}+o(1)}+KL^{2}M\left(1+\frac{MN}{p^{1+\frac{1}{r}}}\right)p^{o(1)}$
@@
$\qquad\qquad+(LM)^{2}p^{\frac{1}{r}+o(1)}+KL^{2}M\left(1+\frac{MN}{p^{1+\frac{1}{r}}}\right)p^{o(1)}.$
@@
Therefore, Theorem 2.11 is proved as soon as we have the inequality
@@
$a=b=1,\,\alpha_{m}={\bf 1}_{{\mathcal{M}}}(m)\text{ and }\beta_{k,n}={\bf 1}_{[1,K]}(k)\cdot{\bf 1}_{{\mathcal{N}}}(n),$
now (7.2) is proved as soon as one has the following four inequalities
$M\geqslant K^{\frac{1}{2}},\quad(MN)^{2}\leqslant Kp^{1+\frac{1}{r}},\quad K(MN)^{2}\geqslant p^{1+\frac{1}{r}+\frac{\eta}{2}},\quad M\geqslant p^{\frac{\eta}{4}}.$
The choice (7.1) guarantees the above four conditions provided that

### v4

@@ 6.2. Bounding $\Sigma_{2}$
We substitute the bounds (6.13), (6.14) and (6.27) into (6.12) and note that
$d_{1}d_{2}=\frac{d^{2}}{\gcd(a,d)\gcd(b,d)}=\frac{d^{2}}{\gcd(ab,d)}$
since we have assumed $\gcd(a,b)=1$. Hence, we derive
$Q_{2}(d)$ $\leqslant p^{o(1)}\sup_{\begin{subarray}{c}C\leqslant K/d_{1}\end{subarray}}\left(CL^{2}M^{2}NUWd_{2}^{-1}\right)^{1-\frac{1}{r}}(M/C+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}\left(\frac{\mathscr{L}_{2}^{2}LM}{dV}\right)^{\frac{1}{2r}}$
×$\qquad\qquad\qquad\qquad\qquad\times\left(CW\gcd(b,d)+NU\gcd(a,d)/d_{2}\right)^{\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}.$
It is easy to see that in the last expression, after expanding, $C$ appears only in positive powers, and therefore the supremum is attained at $C=K/d_{1}$. Hence,
$Q_{2}(d)$ $\leqslant p^{o(1)}\left(KL^{2}M^{2}NUW/d^{2}\right)^{1-\frac{1}{r}}(d_{1}M/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}\left(\frac{\mathscr{L}_{2}^{2}LM}{dV}\right)^{\frac{1}{2r}}$
×$\qquad\times\gcd(ab,d)^{1-\frac{1}{r}}\left(KW\gcd(b,d)/d_{1}+NU\gcd(a,d)/d_{2}\right)^{\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}$
$\leqslant p^{o(1)}\left(KL^{2}M^{2}NUW/d^{2}\right)^{1-\frac{1}{r}}(d_{1}M/K+1)^{\frac{1}{r}}\left(1+\frac{\mathscr{L}_{2}^{2}}{dpV}\right)^{\frac{1}{r}}\left(\frac{\mathscr{L}_{2}^{2}KLMN}{d^{2}V^{2}}\right)^{\frac{1}{2r}}$
×$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\times\gcd(ab,d)^{1-\frac{1}{2r}}(V^{2r}p+V^{r}p^{2})^{\frac{1}{2r}}.$
We also note that
$d^{-2+\frac{2}{r}}\cdot d_{1}^{\frac{1}{r}}\cdot d^{-\frac{1}{r}}\cdot\gcd(ab,d)^{1-\frac{1}{2r}}\leqslant d^{-2+\frac{2}{r}}\gcd(ab,d)^{1-\frac{1}{2r}}\leqslant d^{-1}\gcd(ab,d)^{1-\frac{1}{2r}}$
fo $r\geqslant 2$. Recalling (6.7), (6.9) and (6.11) and then summing over $d\leqslant D$, we obtain
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+KL^{2}M^{2}U^{-1}p^{o(1)}+KL^{2}M\left(V+\frac{MN}{p}\right)p^{o(1)}.$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(LM)^{2}p^{\frac{1}{r}+o(1)}+KL^{2}M\left(p^{\frac{1}{r}}+\frac{MN}{p}\right)p^{o(1)}$
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(LM)^{2}p^{\frac{1}{r}+o(1)}+KL^{2}M\left(p^{\frac{1}{r}}+\frac{MN}{p}\right)p^{o(1)}.$
@@
Therefore, Theorem 2.11 is proved as soon as we have the bound
@@
$a=b=1,\qquad\alpha_{m}={\bf 1}_{{\mathcal{M}}}(m),\qquad\beta_{k,n}={\bf 1}_{[1,K]}(k)\cdot{\bf 1}_{{\mathcal{N}}}(n),$
now (7.2) is proved as soon as one has the following three inequalities
$M\geqslant K^{\frac{1}{2}},\qquad K(MN)^{2}\geqslant p^{1+\frac{1}{r}+\frac{\eta}{2}},\qquad M\geqslant p^{\frac{\eta}{4}}.$
The choice (7.1) guarantees the above three conditions provided that